OpenStudy (anonymous):

Greg combined different weights of fertilizer, soil, and compost to make 29 kilograms (kg) of potting soil for some plants. The weight of fertilizer was x kg. The weight of soil was 2 kg more than three times the weight of fertilizer, and the weight of compost was half the weight of fertilizer. This is represented in the equation below. x + (3x + 2) + 1/2 x = 29 What was the difference in the weights of soil and compost Greg combined?

OpenStudy (anonymous):

4.5x+2 = 29 -2 -2 4.5x = 27 /4.5 /4.5 x = 6 x = weight of fertilizer = 6kg weight of soil = 3x +2= 3(6) +2 = 20kg weight of compost = (1/2)x = 0.5(6) = 3kg